find the two consecutive odd integers such that twice the smaller decreased by the larger equals 53
Answer by checkley71(8403)

LET X & (X+2) BE THE 2 INTEGERS.
2X-(X+2)=53
2X-X-2=53
X=53+2
X=55 ANSWER FOR THE SMALLER ODD NUMBER.
55+2=57 FOR THE LARGER ODD INTEGER.
PROOF
2*55-(55+2)=53
110-57=53
53=53